# Long-Context Linear System Identification

### v1

@@ B.3 Lower and upper isometries
By the properties of Frobenius norm and Lemma B.12,
$\begin{split}\|\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\|_{F}^{2}&=\|(\Delta_{{\boldsymbol{A}}_{1}}-\Delta_{{\boldsymbol{A}}_{2}}+\Delta_{{\boldsymbol{A}}_{2}})L_{\star}E\|_{F}^{2}\\ &\leqslant 2\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}+2\|(\Delta_{{\boldsymbol{A}}_{1}}-\Delta_{{\boldsymbol{A}}_{2}})L_{\star}E\|_{F}^{2}\\ &\leqslant 2\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}+2\|\Delta_{{\boldsymbol{A}}_{1}}-\Delta_{{\boldsymbol{A}}_{2}}\|_{\rm op}^{2}\|L_{\star}\|_{\rm op}^{2}\|E\|_{F}^{2}\\ &\leqslant 2\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}+2p^{\prime}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}^{2}\|L_{\star}\|_{\rm op}^{2}\|E\|_{F}^{2}\,.\end{split}$ (30)
The results readily follows by reordering terms. ∎
Lemma B.23.
@@
$\begin{split}\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E)&\leqslant\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E)+\sqrt{p^{\prime}}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}\|L_{\star}\|_{\rm op}\|E\|_{F}^{2}\,,\\ \end{split}$ (31)
@@
By the properties of trace and Lemma B.12,
$\begin{split}\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E)&=\operatorname{Tr}\left(E^{\top}(\Delta_{{\boldsymbol{A}}_{2}}-\Delta_{{\boldsymbol{A}}_{1}}+\Delta_{{\boldsymbol{A}}_{1}})L_{\star}E\right)\\ &=\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\right)+\operatorname{Tr}\left(E^{\top}(\Delta_{{\boldsymbol{A}}_{2}}-\Delta_{{\boldsymbol{A}}_{1}})L_{\star}E\right)\\ &\leqslant\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\right)+\|\Delta_{{\boldsymbol{A}}_{2}}-\Delta_{{\boldsymbol{A}}_{1}}\|_{\rm op}\|L_{\star}\|_{\rm op}\|E\|_{F}^{2}\\ &\leqslant\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\right)+\sqrt{p^{\prime}}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}\|L_{\star}\|_{\rm op}\|E\|_{F}^{2}\,.\\ \end{split}$ (32)
B.4 Concentration inequalities
In Remark B.24, we show that the quantities of interest that show up in Equation 21 are related to a martingale series and its predictable quadratic variation. This allow us to use Lemma A.5 in Theorem B.26 to quantify the probability of the event in Equation 21 for finite sets of ${\boldsymbol{A}}$.
Remark B.24.
@@ B.4 Concentration inequalities
$d_{t,i}^{(n)}=\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)x_{t}^{(n)}\right)_{i}\left(\xi_{t}^{(n)}\right)_{i}/\sigma^{2}\,,$
@@
Let $W_{\boldsymbol{A}}^{R}$ be the quadratic variation of the series plus an error term as in Theorem A.4, i.e.,
@@
$\begin{split}Y_{\boldsymbol{A}}&=\frac{1}{\sigma^{2}}\sum_{n,t}\langle({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}})X_{t}^{(n)},\xi_{t}^{(n)}\rangle=\frac{1}{\sigma^{2}}\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}}L_{\star}E\right)\,,\\ W_{\boldsymbol{A}}&\coloneqq W_{\boldsymbol{A}}^{0}=\frac{1}{\sigma^{2}}\sum_{n,t}\left\|\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)X_{t}^{(n)}\right\|^{2}=\frac{1}{\sigma^{2}}{{\left\|\Delta_{{\boldsymbol{A}}}L_{\star}E\right\|}}_{F}^{2}\,.\end{split}$
Proposition B.25.
Let $R>0$ be a constant and $\delta>0$ be small. Then, there exist a constant $C^{\prime}(\delta)=\mathcal{O}(\ln\frac{1}{\delta})$ such that
$\forall{\boldsymbol{A}}\in\mathcal{A}(D),\quad C^{\prime}(\delta)\ln dTN\cdot W_{{\boldsymbol{A}}}\geqslant W_{{\boldsymbol{A}}}^{R}\,,$ (33)
with probability $1-\delta$.
@@
By Corollary A.7, there exist a constant $c^{\prime}(\delta)=\mathcal{O}\left(\sqrt{\ln\frac{1}{\delta}}\right)$,
$\sup_{t,n}\|\xi_{t}^{(n)}\|_{\infty}\leqslant c^{\prime}(\delta)\sigma\sqrt{2\ln dTN}\,.$
@@
$\begin{split}W_{{\boldsymbol{A}}}^{R}&=W_{{\boldsymbol{A}}}+\sum_{n,t,i}\mathds{1}_{d_{t,i}^{(n)}>R}\left(d_{t,i}^{(n)}\right)^{2}\\ &\leqslant W_{{\boldsymbol{A}}}+2c^{\prime}(\delta)^{2}\ln dTN\sum_{n,t,i}\mathds{1}_{d_{t,i}^{(n)}>R}\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)x_{t}^{(n)}\right)_{i}^{2}\\ &\leqslant W_{{\boldsymbol{A}}}+2c^{\prime}(\delta)^{2}\ln dTN\sum_{n,t,i}\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)x_{t}^{(n)}\right)_{i}^{2}\\ &\leqslant W_{{\boldsymbol{A}}}+2c^{\prime}(\delta)^{2}\ln dTN\cdot W_{{\boldsymbol{A}}}\,.\end{split}$
Then, by rearrainging terms, we have
$\left(1+2c^{\prime}(\delta)^{2}\ln dTN\right)W_{{\boldsymbol{A}}}\geqslant W_{{\boldsymbol{A}}}^{R}\,.$
Theorem B.26.
Let $\mathcal{S}\subseteq\mathcal{A}(D)$ be a set and let $\mathcal{E}(\mathcal{S})$ be the following event
@@
where $\alpha_{L},\alpha_{U}>0$ are two constants. Then, for any $\gamma>0,R>1$ and $\delta>0$ small, there exist a constant $C^{\prime}(\delta)=\mathcal{O}(\ln\frac{1}{\delta})$ such that
$\begin{split}\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant\gamma Y_{{\boldsymbol{A}}}\right)\leqslant|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{2C^{\prime}(\delta)R\gamma\ln dTN}+\ln\left(\ln\left(\dfrac{\alpha_{U}}{R\alpha_{L}}\right)+1\right)\right)+\mathbb{P}\left(\mathcal{E}(\mathcal{S})^{C}\right)+\delta\,.\end{split}$
@@
The statement is trivial for $\alpha_{L}\leqslant\alpha_{U}$ so we consider the case $\alpha_{L}<\alpha_{U}$.
@@
$\mathcal{E}_{{\boldsymbol{A}}}={{\left\{W_{{\boldsymbol{A}}}^{R}\geqslant\alpha_{L}\right\}}}\cap{{\left\{\gamma Y_{{\boldsymbol{A}}}\leqslant\alpha_{U}\right\}}}\,.$
@@
($\begin{split}\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\right)&=\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)+\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})^{C}\right)\\ &\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)+\mathbb{P}\left(\mathcal{E}(\mathcal{S})^{C}\right)\\ &\leqslant\sum_{{\boldsymbol{A}}\in\mathcal{S}}\mathbb{P}\left({{\left\{W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)+\mathbb{P}\left(\mathcal{E}(\mathcal{S})^{C}\right)\\ &\leqslant\sum_{{\boldsymbol{A}}\in\mathcal{S}}\mathbb{P}\left({{\left\{W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}_{{\boldsymbol{A}}}\right)+\mathbb{P}\left(\mathcal{E}(\mathcal{S})^{C}\right)\,.\end{split}$
@@
$\mathbb{P}\left(W_{\boldsymbol{A}}^{R}\leqslant\gamma Y_{\boldsymbol{A}}\cap\mathcal{E}_{\boldsymbol{A}}\right)\leqslant\exp\left(-\dfrac{\alpha_{L}}{2R\gamma}+\ln\ln\left(\dfrac{\alpha_{U}}{R\alpha_{L}}\right)\right)\,.$
by Lemma A.5 which implies that
@@
($\begin{split}\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\right)\leqslant|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{2R\gamma}+\ln\left(\ln\left(\dfrac{\alpha_{U}}{R\alpha_{L}}\right)+1\right)\right)+\mathbb{P}\left(\mathcal{E}(\mathcal{S})^{C}\right)\,.\end{split}$
Finally, let $C^{\prime}(\delta)$ be the constant from Proposition B.25 and $\mathcal{E}_{\delta}$ be the event in Equation 33. Then,
@@
($\begin{split}\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\right)&\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}_{\delta}\right)+\mathbb{P}\left(\mathcal{E}_{\delta}^{C}\right)\\ &\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant C^{\prime}(\delta)\gamma\ln dTN\cdot Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}_{\delta}\right)+\delta\\ &\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant C^{\prime}(\delta)\gamma\ln dTN\cdot Y_{{\boldsymbol{A}}}\right\}}}\right)+\delta\,.\end{split}$
B.5 Proof of Theorem B.6
By Corollary B.17, there exist a constant $C(\delta/4)$ such that for all $C\geqslant C(\delta/4)$
@@
$\begin{split}\inf_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}\,,\\ \sup_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\frac{\sigma^{2}}{2}\|L_{\star}\|_{\rm op}C\eta Dp^{\prime}dr\left(\ln\tau\right)^{2}\,,\end{split}$ (34)
with probability $41-\delta/$ 4.
Let $\mathcal{S}$ be an $\epsilon$-net over $\mathcal{A}(D)\setminus\mathcal{G}(C,D)$. Then, by Theorem B.26, there exist a constant $C^{\prime}(\delta/4)$ such that
@@
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{S}:\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant\gamma\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\leqslant\\ &\quad\quad\quad\quad|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{2C^{\prime}(\delta/4)R\gamma\ln dTN}+\ln\left(\ln\left(\dfrac{\alpha_{U}}{R\alpha_{L}}\right)+1\right)\right)+\frac{2\delta}{4}\,,\end{split}$
where $\alpha_{L}$ and $\alpha_{U}$ are the lower and upper bounds in Equation 34 scaled with $\sigma^{2}$:
@@
$\begin{split}\alpha_{L}&=\frac{1}{8}\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}\,,\\ \alpha_{U}&=\frac{1}{2}\|L_{\star}\|_{\rm op}C\eta Dp^{\prime}dr\left(\ln\tau\right)^{3/2}\,.\end{split}$

### v2

@@ C.3 Lower and upper isometries
By the properties of Frobenius norm and Lemma C.4,
$\begin{split}\|\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\|_{F}^{2}&=\|(\Delta_{{\boldsymbol{A}}_{1}}-\Delta_{{\boldsymbol{A}}_{2}}+\Delta_{{\boldsymbol{A}}_{2}})L_{\star}E\|_{F}^{2}\\ &\leqslant 2\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}+2\|(\Delta_{{\boldsymbol{A}}_{1}}-\Delta_{{\boldsymbol{A}}_{2}})L_{\star}E\|_{F}^{2}\\ &\leqslant 2\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}+2\|\Delta_{{\boldsymbol{A}}_{1}}-\Delta_{{\boldsymbol{A}}_{2}}\|_{\rm op}^{2}\|L_{\star}\|_{\rm op}^{2}\|E\|_{F}^{2}\\ &\leqslant 2\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}+2p^{\prime}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}^{2}\|L_{\star}\|_{\rm op}^{2}\|E\|_{F}^{2}\,.\end{split}$
The result readily follows by reordering terms. ∎
Lemma C.9.
@@
$\begin{split}\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E)&\leqslant\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E)+\sqrt{p^{\prime}}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}\|L_{\star}\|_{\rm op}\|E\|_{F}^{2}\,,\\ \end{split}$
@@
By the properties of trace and Lemma C.4,
$\begin{split}\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E)&=\operatorname{Tr}\left(E^{\top}(\Delta_{{\boldsymbol{A}}_{2}}-\Delta_{{\boldsymbol{A}}_{1}}+\Delta_{{\boldsymbol{A}}_{1}})L_{\star}E\right)\\ &=\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\right)+\operatorname{Tr}\left(E^{\top}(\Delta_{{\boldsymbol{A}}_{2}}-\Delta_{{\boldsymbol{A}}_{1}})L_{\star}E\right)\\ &\leqslant\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\right)+\|\Delta_{{\boldsymbol{A}}_{2}}-\Delta_{{\boldsymbol{A}}_{1}}\|_{\rm op}\|L_{\star}\|_{\rm op}\|E\|_{F}^{2}\\ &\leqslant\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\right)+\sqrt{p^{\prime}}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}\|L_{\star}\|_{\rm op}\|E\|_{F}^{2}\,.\\ \end{split}$
C.4 Concentration inequalities
In Remark C.2, we show that the quantities of interest that show up in Equation 19 are related to a martingale series and its predictable quadratic variation. This allow us to use Lemma B.1 in Theorem C.4 to quantify the probability of the event in Equation 19 for finite sets of ${\boldsymbol{A}}$.
Remark C.2.
@@ C.4 Concentration inequalities
$d_{t,i}^{(n)}=\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)X_{t}^{(n)}\right)_{i}\left(\xi_{t}^{(n)}\right)_{i}\,,$
@@
Let $W_{\boldsymbol{A}}^{R}$ be the quadratic variation of the series plus an error term as in Theorem B.3, i.e.,
@@
$\begin{split}Y_{\boldsymbol{A}}&=\sum_{n,t}\langle({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}})X_{t}^{(n)},\xi_{t}^{(n)}\rangle=\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}}L_{\star}E\right)\,,\\ W_{\boldsymbol{A}}&\coloneqq W_{\boldsymbol{A}}^{0}=\sigma^{2}\sum_{n,t}\left\|\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)X_{t}^{(n)}\right\|^{2}=\sigma^{2}{{\left\|\Delta_{{\boldsymbol{A}}}L_{\star}E\right\|}}_{F}^{2}\,.\end{split}$
Proposition C.3.
Let $0<\delta<e^{-1}$ and $R>0$ be constants. Then, with probability at least $1-\delta$, we have
$\forall{\boldsymbol{A}}\in\mathcal{A}(D),\quad C^{\prime}\left(\ln 2dNT+\ln\dfrac{1}{\delta}\right)W_{{\boldsymbol{A}}}\geqslant W_{{\boldsymbol{A}}}^{R}\,,$ (25)
where $C^{\prime}=1+4c^{\prime 2}$ and $c^{\prime}$ is the universal constant in Lemma B.2.
@@
By Corollary B.1, there exists a constant $c^{\prime}$ such that
$\sup_{t,n}\|\xi_{t}^{(n)}\|_{\infty}\leqslant c^{\prime}\sigma\sqrt{2}\left(\sqrt{\ln 2dTN}+\sqrt{\ln\dfrac{1}{\delta}}\right)\,.$
@@
$\begin{split}W_{{\boldsymbol{A}}}^{R}&=W_{{\boldsymbol{A}}}+\sum_{n,t,i}\mathds{1}_{d_{t,i}^{(n)}>R}\left(d_{t,i}^{(n)}\right)^{2}\\ &\leqslant W_{{\boldsymbol{A}}}+4c^{\prime 2}\sigma^{2}\left(\ln 2dTN+\ln\dfrac{1}{\delta}\right)\sum_{n,t,i}\mathds{1}_{d_{t,i}^{(n)}>R}\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)X_{t}^{(n)}\right)_{i}^{2}\\ &\leqslant W_{{\boldsymbol{A}}}+4c^{\prime 2}\sigma^{2}\left(\ln 2dTN+\ln\dfrac{1}{\delta}\right)\sum_{n,t,i}\left(\left({\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right)X_{t}^{(n)}\right)_{i}^{2}\\ &\leqslant W_{{\boldsymbol{A}}}+4c^{\prime 2}\left(\ln 2dTN+\ln\dfrac{1}{\delta}\right)W_{{\boldsymbol{A}}}\,.\end{split}$
Then, by rearranging terms, we have
$\left(1+4c^{\prime 2}\left(\ln 2dTN+\ln\dfrac{1}{\delta}\right)\right)W_{{\boldsymbol{A}}}\geqslant W_{{\boldsymbol{A}}}^{R}\,.$
Theorem C.4.
Let $\mathcal{S}\subseteq\mathcal{A}(D)$ be a finite set and let $\mathcal{E}(\mathcal{S})$ be the following event
@@
where $\alpha_{L},\alpha_{U}>0$ are two constants. Then, for any $\gamma,R>0$ and $0<\delta<e^{-1}$,
($\begin{split}&\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)\\ &\quad\leqslant|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{2\gamma^{\prime}\left(\gamma^{\prime}+R\right)}+\ln\left(\ln\left(\dfrac{\gamma^{\prime}\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)+\delta\,,\end{split}$
where $\gamma^{\prime}=C^{\prime}e\gamma\left(\ln 2dTN+\ln\dfrac{1}{\delta}\right)$ and $C^{\prime}$ is the universal constant in Proposition C.3.
@@
$\mathcal{E}_{{\boldsymbol{A}}}={{\left\{W_{{\boldsymbol{A}}}^{R}\geqslant\alpha_{L}\right\}}}\cap{{\left\{Y_{{\boldsymbol{A}}}\leqslant\alpha_{U}\right\}}}\,.$
@@
($\begin{split}\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)&\leqslant\sum_{{\boldsymbol{A}}\in\mathcal{S}}\mathbb{P}\left({{\left\{W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)\\ &\leqslant\sum_{{\boldsymbol{A}}\in\mathcal{S}}\mathbb{P}\left({{\left\{W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}_{{\boldsymbol{A}}}\right)\,.\end{split}$
@@
$\mathbb{P}\left(W_{\boldsymbol{A}}^{R}\leqslant\gamma Y_{\boldsymbol{A}}\cap\mathcal{E}_{\boldsymbol{A}}\right)\leqslant\exp\left(-\dfrac{\alpha_{L}}{2e\gamma\left(e\gamma+R\right)}+\ln\left(\ln\left(\dfrac{\gamma\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)\,.$
by Lemma B.1 which implies that
@@
($\begin{split}\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)\leqslant|\mathcal{S}|\exp\left(-\dfrac{\alpha_{L}}{2e\gamma\left(e\gamma+R\right)}+\ln\left(\ln\left(\dfrac{\gamma\alpha_{U}}{\alpha_{L}}\right)+1\right)\right)\,.\end{split}$
Finally, let $\mathcal{E}_{\delta}$ be the event in Equation 25. Then,
@@
($\begin{split}&\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)\\ &\quad\quad\quad\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}\leqslant\gamma Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\cap\mathcal{E}_{\delta}\right)+\mathbb{P}\left(\mathcal{E}_{\delta}^{C}\right)\\ &\quad\quad\quad\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant C^{\prime}\gamma\left(\ln 2dNT+\ln\dfrac{1}{\delta}\right)Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\cap\mathcal{E}_{\delta}\right)+\delta\\ &\quad\quad\quad\leqslant\mathbb{P}\left({{\left\{\exists{\boldsymbol{A}}\in\mathcal{S}:W_{{\boldsymbol{A}}}^{R}\leqslant C^{\prime}\gamma\left(\ln 2dNT+\ln\dfrac{1}{\delta}\right)Y_{{\boldsymbol{A}}}\right\}}}\cap\mathcal{E}(\mathcal{S})\right)+\delta\,.\end{split}$
C.5 Proof of Theorem C.2
Recall that Remark C.2 shows that
@@
$\begin{split}Y_{\boldsymbol{A}}=\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{A}}}L_{\star}E\right)\,,\quad W_{\boldsymbol{A}}=\sigma^{2}{{\left\|\Delta_{{\boldsymbol{A}}}L_{\star}E\right\|}}_{F}^{2}\,.\end{split}$
Therefore, we need to show that for any $0<\delta<e^{-1}$, there exists a constant $C(\delta)$ such that
@@
$\begin{split}\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C(\delta),D):W_{\boldsymbol{A}}\leqslant 2\sigma^{2}\eta Y_{\boldsymbol{A}}\right)\leqslant\delta\,.\end{split}$
By Corollary C.5, there exists a constant $C_{1}(\delta/4)=\mathcal{O}(\ln(1/\delta))$ such that for all $C\geqslant C_{1}(\delta/4)$
@@
$\begin{split}\inf_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}W_{{\boldsymbol{A}}}&\geqslant\frac{\sigma^{4}}{8}\sigma_{\min}(L_{\star})^{2}CD^{2}\eta^{2}\tau p^{\prime}dr\,,\\ \sup_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}Y_{{\boldsymbol{A}}}&\leqslant\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{CD^{2}\eta^{2}\tau p^{\prime}d^{2}rNT}\,,\end{split}$ (26)
with probability $41-\delta/$ 4. In addition, by Lemma C.7, with probability at least $41-\delta/$ 4, we have
$\|E\|_{F}^{2}\leqslant\left(1+c^{2}\nu(\delta/4)\right)dNT\,,\quad\text{where}\quad\nu(\delta/4)=\max{{\left\{1,\dfrac{\ln\frac{\delta}{4}}{C_{HW}dNT}\right\}}}\,.$ (27)
In the following, we work conditionally on these events.
Let $\mathcal{S}$ be an $\epsilon$-net over $\mathcal{A}(D)\setminus\mathcal{G}(C,D)$. By Lemmas C.8 and C.9, for any ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$, there exists ${\boldsymbol{A}}^{\prime}\in\mathcal{S}$ such that
$\begin{split}W_{{\boldsymbol{A}}}&\geqslant\frac{1}{2}W_{{\boldsymbol{A}}^{\prime}}-\epsilon_{1}(\delta/4)\epsilon^{2}\,,\quad Y_{{\boldsymbol{A}}}\leqslant Y_{{\boldsymbol{A}}^{\prime}}+\epsilon_{2}(\delta/4)\epsilon\,,\\ \end{split}$
where $\epsilon_{1}(\delta/4),\epsilon_{2}(\delta/4)$ are given as follows:
$\begin{split}\epsilon_{1}(\delta/4)&=\sigma^{4}(1+c^{2}\nu_{3}(\delta/4))p^{\prime}dNT\|L_{\star}\|_{\rm op}^{2}\,,\\ \epsilon_{2}(\delta/4)&=\sigma^{2}(1+c^{2}\nu_{3}(\delta/4))\sqrt{p^{\prime}}dNT\|L_{\star}\|_{\rm op}\,.\end{split}$
We set $\epsilon$ as follows:
$\epsilon=\min{{\left\{\sqrt{\dfrac{\alpha_{L}}{\epsilon_{1}(\delta/4)}},\dfrac{\alpha_{L}}{4\sigma^{2}\eta\epsilon_{2}(\delta/4)},D\right\}}}\,.$
In particular, $\epsilon$ is small such that for any ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$,
$\exists{\boldsymbol{A}}^{\prime}\in\mathcal{S}:\quad W_{{\boldsymbol{A}}^{\prime}}\leqslant 2W_{{\boldsymbol{A}}}+2\alpha_{L}\,,\quad Y_{{\boldsymbol{A}}^{\prime}}\geqslant Y_{{\boldsymbol{A}}}-\frac{\alpha_{L}}{4\sigma^{2}\eta}\,.$ (28)
Assume that ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$ verifies
$W_{{\boldsymbol{A}}}\leqslant 2\sigma^{2}\eta Y_{{\boldsymbol{A}}}\,.$
Then, by Equation 28, there exists ${\boldsymbol{A}}^{\prime}\in\mathcal{S}$ such that
$W_{{\boldsymbol{A}}^{\prime}}\leqslant 4W_{{\boldsymbol{A}}}\leqslant 8\sigma^{2}\eta Y_{{\boldsymbol{A}}}\leqslant 16\sigma^{2}\eta Y_{{\boldsymbol{A}}}-4W_{{\boldsymbol{A}}}\leqslant 16\sigma^{2}\eta Y_{{\boldsymbol{A}}^{\prime}}\,.$